I cant believe how bad my brother is at growing things. He bought a whole bunch of plants at the nursery the other day, right away five died. then our dog dug up 2/9 of them. out of the remaining plants half of them died pretty soon afterword then a rabbit came and ate half of what was left and half a plant more. my brother only had three plants left then how many did he buy at the nursery?
This is what I have so far:5 died 2/9 dug up (so 7/9 left) 1/2 died (so 1/2 left) 1/2 eaten (so 1/2 left) and 1/2 a plant eaten
can you please write the solution out step by step

Found 2 solutions by stanbon, solver91311:
Answer by stanbon(75887)

I cant believe how bad my brother is at growing things. He bought a whole bunch of plants at the nursery the other day, right away five died. then our dog dug up 2/9 of them. out of the remaining plants half of them died pretty soon afterword then a rabbit came and ate half of what was left and half a plant more. my brother only had three plants left then how many did he buy at the nursery?
This is what I ha]ve so far:5 died 2/9 dug up (so 7/9 left) 1/2 died (so 1/2 left) 1/2 eaten (so 1/2 left) and 1/2 a plant eaten
-----
Bought:: x plants
Equation:
x-5
(7/9)(x-5)
(1/2)[(7/9)(x-5)]
(1/2)^2[(7/9)(x-5)]-(1/2) = 3
----------------------------------
(1/4)(7/9)(x-5)] = 3.5
-----
(7/36)(x-5) = 7/2
-----
(1/36)(x-5) = 1/2
x-5 = 18
Ans: x = 23
----------------
Cheers,
Stan H.
----------

Answer by solver91311(24713)



Work backwards.

3 at the end.

Let be the number of plants that were there just before the rabbit began eating. So there were of which the rabbit ate one-half plus one-half of a plant, namely leaving 3. Putting it together, just before the rabbit arrived:







So just before the rabbit there were 7 plants.

Then, right after the dog, just before half of them died, there were twice as many, namely 14.

If the dog dug up , then what was left (the 14) must have been of what was there before the dog got there, so let that amount be and we have:





Adding in the 5 that died right away, you must have started with 23.

John

My calculator said it, I believe it, that settles it